Use the direct stiffness method and calculate the forces, deflections and moments in each node.

Mass of UDL = 115 kg

E=200 Gpa

I= 4 x 10^-5 m^4

The downward arrow's force = 1 128,15 N

UDL wa O ह ह 14 Ri Rz

.
Answer #1

F=1128 VOL RI R2 F = Given up = 115bg, 1150N 19.17 N/m. By Equilibrium conditions. F M Efy co. R, & R₂ - 1150 & 11 28.15 RitRconsider section Box V= 1139.1 19.12 x aloexes m(x) = 0.15 + 1139.1 x for section (Be) v= 1 139.1 - 19.178 1128.15 m.r) = 0.1